OpenStudy (anonymous):

The side of a triangle is 10 square meters longer than the side of an equalateral triangle. The perimeter of the square is 3 times the perimeter of the triangle. Find the length of each side of the triangle.

OpenStudy (tcp2315):

the side is measured in square meters?

OpenStudy (anonymous):

no, it just says meters

OpenStudy (tcp2315):

re read the first line of your question, is everything correctly typed? just curious if you mean the side of a square is 10 meters longer then the side of a equilateral triangle

OpenStudy (anonymous):

A side of a square is 10 meters longer than the side of an equalateral triangle.

OpenStudy (tcp2315):

ok, let x= length of the square and y=length of triangle so x=y+10 and 4x=3(3y) 4x=9y rearragne the first equation for y, and sub it in 4x=9(x-10) 4x=9x-90 -5x=-90 x=18 meters :)